# NoiseNCA: Noisy Seed Improves Spatio-Temporal Continuity of Neural Cellular Automata

### v2

@@ Introduction
To find the fixed point state of an NCA, we numerically solve the optimization problem in Equation 8 using gradient descent. To visualize the obtained fixed point, we set the state of all cells to be equal to this fixed point state and show the results in the third row of Figure9 . These results show that for textures in Group A, applying the NCA update rule with small $\Delta\textup{t}$ (illustrated in the second row of Figure 9) converges to the same state as the fixed point solution given by Equation 8. However, for textures in Group B, the model synthesizes the correct pattern using small values for $\Delta\textup{t}$, showing that the NCA output does not converge to the fixed point solution. We hypothesize that this behavioral difference between Groups A and B is due to the stability of the fixed point state. In practice, the value of $\Delta\mathrm{S}_{fp}$ is very close to but not exactly zero; therefore, the NCA model can escape from an unstable or saddle fixed point. We find that if we initialize the cell state with $\mathrm{S}_{fp}$ and apply the NCA update rule for a large number of steps, the NCA will escape the fixed point state and converge to the correct pattern for textures in Group B, while it will remain in the fixed point state for textures in Group A. This shows that the fixed point state is stable for textures in Group A and is a saddle point for textures in Group B. We speculate that for textures in Group A, when $\Delta\textup{t}$ is small, the stochastic update schema does not create enough variation between cells, which causes the NCA to converge to a fixed point with uniform state.
@@
The cell size given by the values $\Delta\textup{x},\Delta\textup{y}$ in Equation 5 controls the scale of the patterns. The smaller the cell size, the more cells (pixels) are required to create the pattern; therefore, the pattern becomes larger. In this section, we demonstrate three interesting ways to control and exploit the scale of the patterns at inference time.
@@
By setting $\Delta\textup{x},\Delta\textup{y}$ to be functions of the cell coordinate $x,y$, we can have local control over the scale of the synthesized patterns. In this example, we use a grid of size $256\times 1536$ and vary the cell size as a function of the $x$ coordinate, that is, $\Delta\textup{x}=\Delta\textup{y}=g(x)$. Figure 10 shows the results for 4 different textures. The target texture is shown on the left and the synthesized images demonstrating multiscale patterns are shown on the right. The function $g(x)$ is symmetric around the center of the image, increasing exponentially from $2^{-3.0}$ on the left border to $2^{0.5}$. This makes the patterns on the border to be roughly 11.0 times larger than the patterns in the center. Note how the scale of the patterns change smoothly and seamlessly across the image.
@@
Noise-NCA can create Anisotropically scaled patterns when $\Delta\textup{x}\neq\Delta\textup{y}$. In this example, we use two different values for each dimension of the cell size $\Delta\textup{x},\Delta\textup{y}\in\left\{1.0,0.5\right\}\times\left\{1.0,0.5\right\}$ and set the grid size to $128\times 128$ regardless of the cell size. Figure 11 shows the anisotropic patterns for six different textures. Decreasing $\Delta\textup{x}$ is like stretching patterns horizontally, while decreasing $\Delta\textup{y}$ acts as a vertical stretch.
@@
Varying the cell size $\Delta\textup{x},\Delta\textup{y}$ over time $t$ causes the patterns to grow or shrink. We find that if the scale variation is smooth enough, the patterns remain coherent. For example, for the bubble pattern shown in Figure 12, when we decrease $\Delta\textup{x},\Delta\textup{y}$ over time, some bubbles grow coherently in size without losing the whole bubble structure, as indicated by the red circle.
@@ Conclusion
In this paper, we investigate whether NCA learns a continuous PDE that governs its dynamics or simply overfits the discretization used during training, by studying the NCA behavior at the continuous space-time limit. We show that existing NCA models struggle to generate correct textures when we change the space or time discretization, indicating overfitting. We analyze the source of the overfitting problem and propose Noise-NCA, an effective and easy-to-implement solution based on using random uniform noise as the seed. This change significantly improves NCA’s ability to maintain consistent behavior across different discretization granularities and enables continuous control over the speed of the pattern formation and the scale of the synthesized patterns.

### v3

@@ Introduction
Note that the optimization objective in Equation 8 is convex, and hence all local minimas are connected and every local minima is also a global minima. To find a fixed point state of an NCA, we numerically solve this optimization problem using gradient descent. We find that, regardless of the starting point, the optimization converges nearly to the same solution. This suggests that the fixed point states, i.e. the global minimas, are tightly packed. To visualize the obtained fixed point, we set the state of all cells to be equal to this fixed point state and show the results in the third row of Figure 9 . These results show that for textures in Group A, applying the NCA update rule with small $\Delta\textup{t}$ (illustrated in the second row of Figure 9) converges to the same state as the fixed point solution given by Equation 8. However, for textures in Group B, the model synthesizes the correct pattern using small $\Delta\textup{t}$ values, showing that the NCA output does not converge to the fixed point solution. We hypothesize that this behavioral difference between Groups A and B is due to the stability of the fixed point state. In practice, the value of $\Delta\mathrm{S}_{fp}$ is very close to but not exactly zero; therefore, the NCA model can escape from an unstable or saddle fixed point. We find that if we initialize the cell state with $\mathrm{S}_{fp}$ and apply the NCA update rule for a large number of steps, the NCA will escape the fixed point state and converge to the correct pattern for textures in Group B, while it will remain in the fixed point state for textures in Group A. This shows that the fixed point state is stable for textures in Group A and is a saddle point for textures in Group B. We speculate that for textures in Group A, when $\Delta\textup{t}$ is small, the stochastic update does not create enough variation between cells, which causes the NCA to converge to a uniform state.
@@
The cell size given by the values $(\Delta\textup{x},\Delta\textup{y})$ in Equation 5 controls the scale of the patterns. The smaller the cell size, the more cells (pixels) are required to create the pattern; therefore, the pattern becomes larger. In this section, we demonstrate three interesting ways to control and exploit the scale of the patterns at inference time.
@@
By setting $\Delta\textup{x},\Delta\textup{y}$ to be functions of the cell coordinate $x,y$, we can have local control over the scale of the synthesized patterns. In this example, we use a grid of size $256\times 1536$ and vary the cell size as a function of the $x$ coordinate, that is, $\Delta\textup{x}=\Delta\textup{y}=g(x)$. Figure 10 shows the results for three different textures. The target texture is shown on the left and the synthesized images demonstrating multiscale patterns are shown on the right. The function $g(x)$ is symmetric around the center of the image, increasing exponentially from $2^{-3.0}$ on the left border to $2^{0.5}$. This makes the patterns on the border to be roughly 11.0 times larger than the patterns in the center. Note how the scale of the patterns change smoothly and seamlessly across the image.
@@
Noise-NCA can create anisotropically scaled patterns when $\Delta\textup{x}\neq\Delta\textup{y}$. In this example, we use two different values for each dimension of the cell size $\Delta\textup{x},\Delta\textup{y}\in\left\{1.0,0.5\right\}\times\left\{1.0,0.5\right\}$ and set the grid size to $128\times 128$ regardless of the cell size. Figure 11 shows the anisotropic patterns for six different textures. Decreasing $\Delta\textup{x}$ stretches the patterns horizontally, while decreasing $\Delta\textup{y}$ acts as a vertical stretch.
@@
Varying the cell size $\Delta\textup{x},\Delta\textup{y}$ over time $t$ causes the patterns to grow or shrink. We find that if the scale varies smoothly, the patterns remain coherent. For example, for the bubble pattern shown in Figure 12, when we decrease $\Delta\textup{x},\Delta\textup{y}$ over time, some bubbles grow coherently in size and preserve their soliton-like structure, as indicated by the red circles.
@@ Conclusion
In this paper, by studying the NCA behavior at the continuous space-time limit, we investigate whether NCA learns a continuous PDE that governs its dynamics or simply overfits the discretization used during training. We show that existing NCA models struggle to generate correct textures when we change the space or time discretization, indicating overfitting. We analyze the source of the overfitting problem and propose Noise-NCA, an effective and easy-to-implement solution based on using random uniform noise as the seed. This change significantly improves NCA’s ability to maintain consistent behavior across different discretization granularities and enables continuous control over the speed of the pattern formation and the scale of the synthesized patterns.